# Robust and Flexible Omnidirectional Depth Estimation with Multiple 360 Cameras

### v1

@@ Abstract
Omnidirectional depth estimation has received much attention from researchers in recent years. However, challenges arise due to camera soiling and variations in camera layouts, affecting the robustness and flexibility of the algorithm. In this paper, we use the geometric constraints and redundant information of multiple 360 cameras to achieve robust and flexible multi-view omnidirectional depth estimation. We implement two algorithms, in which the two-stage algorithm obtains initial depth maps by pairwise stereo matching of multiple cameras and fuses the multiple depth maps to achieve the final depth estimation; the one-stage algorithm adopts spherical sweeping based on hypothetical depths to construct a uniform spherical matching cost of the multi-camera images and obtain the depth. Additionally, a generalized epipolar equirectangular projection is introduced to simplify the spherical epipolar constraints. To overcome panorama distortion, a spherical feature extractor is implemented. Furthermore, a synthetic 360 dataset consisting of 12K road scene panoramas and 3K ground truth depth maps is presented to train and evaluate 360 depth estimation algorithms. Our dataset takes soiled camera lenses and glare into consideration, which is more consistent with the real-world environment. Experiments show that our two algorithms achieve state-of-the-art performance, accurately predicting depth maps even when provided with soiled panorama inputs. The flexibility of the algorithms is experimentally validated in terms of camera layouts and numbers.
Index Terms:
Omnidirectional Depth Estimation, Omnidirectional Stereo Matching, Spherical Feature Learning, 360Cameras, Autonomous Driving
I Introduction
Recently, the proliferation of 360 cameras has significantly propelled the advancement of omnidirectional image processing and omnidirectional vision tasks. Omnidirectional depth estimation has attracted attention in numerous applications including autonomous driving and robot navigation, owing to its efficient 3D perception of the 360 environment. Various algorithms have been proposed to estimate omnidirectional depth maps, including monocular [1, 2, 3], binocular [4, 5] and multi-view approaches [6, 7, 8, 9].
The complex geometric constraints and image distortions of spherical images pose challenges for omnidirectional depth estimation. In addition, the camera may be soiled resulting in image degradation in practical applications. As shown in Figure 1(e), the images can be soiled by mud spots, water drops or dazzled by intense light. Apart from this, the camera layouts may vary to accommodate different types of vehicles or robots in real-world tasks. Consequently, the development of an omnidirectional depth estimation algorithm that exhibits robustness against camera soiling and flexibility in adapting to diverse camera configurations becomes imperative and indispensable.
@@ I Introduction
Compared to our conference version[11], this extended work encompasses following advancements. Firstly, we expand the applicability of the Cassini projection, originally designed for left-right 360 camera pairs, to the Generalized Epipolar Equirectangular projection, which accommodates camera pairs at arbitrary relative positions. We also provide a thorough analysis and comparison of the spherical geometry constraint models. We introduce the one-stage Spherical-Sweeping MODE and extensively compare its performance with the two-stage Pairwise Stereo matching methods through a wealth of experiments. Furthermore, we demonstrate the flexibility of the proposed methods with varying layouts and numbers of input cameras. Lastly, we present a comprehensive comparative analysis, encompassing the latest state-of-the-art methods, and provide insights for the future advancement of the field.
II Related Work
II-A Stereo Matching and Multi-view Stereo Methods
@@
Conventional stereo matching methods estimate disparity map based on the stereo epipolar constraint and image features matching. Some methods aggregate global features to achieve high accuracy, such as SGM[12] and its variants[13, 14, 15], and graph-cut based methods[16, 17]. Deep learning methods report much improved performance in stereo matching. Zbontar and Lecun propose MCCNN[18] that implements the feature extraction with CNNs and computes disparity via conventional cost aggregation. Many methods[19, 20, 21] construct 3D cost volume with image features and optimize the 3D-CNN based cost aggeration modules to estimate disparity maps. Some approaches[22, 23] compute the 2D left-right feature correlation volume. AANet[24] adopts an adaptive aggregation algorithm and replaces the costly 3D-CNNs for an efficient architecture. EdgeStereo[25] incorporates the edge cues to improve the quality of disparity in texture-less regions and detailed structures. CFNet[26] proposes a network based on the cascade and fused cost volume to improve the robustness in stereo matching. ACVNet[27] proposes a novel attention-based cost volume construction method to adjust weights of cost volume. AdaStereo[28] proposes multiple level alignment to improve the domain adaptation ability of deep learning models. DMCA-Net[29] utilizes differentiable Markov Random Field for cost aggeration to guide stereo matching. RAFT-Stereo[30] adopts multi-level Gated Recurrent Unit (GRU) to estimate disparity maps recurrently. CREStereo[31] designs a hierarchical network to update disparities iteratively and proposes an adaptive group correlation layer to match points via the local feature.
@@
Multi-view Stereo (MVS) has important applications in 3D reconstruction and has developed rapidly in recent years. Yao et al. [10] proposed the end-to-end MVSNet that builds cost volume by warping feature maps of different views into front-parallel planes of the reference camera to obtain depth maps. P-MVSNet[32] proposes a patch-wise aggregation to build confidence volume and a hybrid network of isotropic and anisotropic 3D-CNNs to exploit context information. Point-MVSNet[33] adopts the feature augmented point cloud to refine the depth map iteratively. Cascade-MVS[34] and CVP-MVS[35] improve the performance with multi-scale coarse-to-fine architectures. UGNet[36] also adopts a coarse-to-fine architecture and leverages uncertainty to improve the depth accuracy. DS-Depth[37] builds the fusion cost volume from multi-frame images to estimate accurate depth maps. PVA-MVSNet[38] proposes self-adaptive view aggregation to generate cost volume instead of the widely-used mean square variance. PVSNet[39] and Vis-MVSNet[40] take the visibility of each view into consideration to supress the mis-matching. Many approaches use the iterative optimization modules to replace the 3DCNNs. R-MVSNet[41] and CER-MVS[42] adopt the GRU module and D2HC-RMVSNet[43] leverages the LSTM module for the cost aggregation.
@@
II-B Omnidirectional Depth Estimation
@@
There are also some methods for obtaining omnidirectional depth maps based on multi-view fisheye cameras. Won et al. propose SweepNet[6] which builds cost volume via spherical sweeping and estimates spherical depth by cost aggregation. They further improve the algorithm and propose the end-to-end OmniMVS[7, 8] architecture to achieve better performance. Meuleman et al. [52] propose an adaptive spherical matching method and an efficient cost aggregation method to achieve real-time omnidirectional MVS. OmniVidar[53] adopts the triple sphere camera model and rectifies the multiple fisheye images into stereo pairs of four directions to obtain depth maps. Su et al.[9] leverage a cascade architecture for cost regularization to achieve high accuracy for omnidirectional detph extimation from four fisheye cameras. However, these methods also use planar convolution to extract spherical features and the blind areas of fisheye cameras introduce discontinuity in the spherical cost volume.
II-C Omnidirectional Depth Datasets
Large-scale datasets with high variety are essential for training and evaluating learning-based algorithms. Recently released omnidirectional depth datasets can be divided into two categories according to the input images, one with the panoramas, and the other with the fisheye images. These datasets are mainly collected from publicly available real-world and synthetic 3D datasets by repurposing them to omnidirectional by rendering. For datasets with panoramas, Wang et al.[48] collect an indoor monocular 360 video dataset named PanoSUNCG from[54]. De La Garanderie et al.[55] provide an outdoor monocular 360 benchmark with 200 images generated from the CARLA autonomous driving simulator[56]. MP3D and SF3D[5] are indoor binocular 360 datasets collected from[57, 58]. 3D60 by Zioulis et al.[47] is an indoor trinocular (central, right, up) 360 dataset collected from[57, 58, 54, 59]. For datasets with fisheye images, Won et al.[6, 7, 8] present three datasets: Urban, OmniHouse and OmniThings. All three datasets are virtually collected in Blender with four fisheye cameras. The fisheye images need complementary information to estimate an omnidirectional depth map, which means discontinuity and requirements for camera directions. In contrast, the panoramas record all 360 information continuously without blind areas. However, as summarized above, the datasets with stereo panoramas consist of indoor scenes only. A detailed summary of multi-view omnidirectional depth datasets can be found in Table I.
III Spherical Geometry Constraint Model
@@
To achieve the robust and accurate depth estimation, we establish the geometry constrint of multiple 360 cameras. In this paper, we introduce two spherical geometry constraint models to leverage the multi-view information. In Section III-A, we introduce the generalized epipolar equirectangular projection, which simplifies the epipolar constraint for binocular panoramas and enables the stereo matching methods on spherical images. In Section III-B, we present the pipeline of spherical sweeping that builds the cost volume of multi-view panoramas based on the hypothetical sphericals.
III-A Generalized Epipolar Equirectangular Projection
@@
In this paper, we propose Generalized epipolar equirectangular (GEER) projection to achieve the epipolar constraint for binocular panoramic cameras at arbitrary relative positions in space. As shown in Figure 2(a), $O_{l}$ and $O_{r}$ are the optic centers of two omnidirectional cameras. We establish a 3D Cartesian coordinate system, where the direction of the x-axis is $O_{r}O_{l}$. $P$ is an object point in 3D space and imaged at points $P_{l}$ and $P_{r}$ on the left and right imaging spheres respectively. $P^{\prime}$ is the projection of $P$ on the plane $yOz$. We define the spherical coordinate system $(\rho,\phi,\theta)$ as follows: $\rho$ is the distance between the object point $P$ and the optic center $O$, $\phi$ is the angle between $PO$ and $x$ axis ($\angle POx$) and denotes the elevation angle, $\theta$ is the angle between $P^{\prime}O$ and $z$ axis on the plane $yOz$ ($\angle P^{\prime}Oz$) and denotes the azimuth angle. Thus, the transformation between Cartesian coordinates and the spherical coordinates is:
@@
where $(u,v)$ denotes the image pixel coordinates in GEER projection and $H,W$ denote the height and width of the image. Because $\theta(\angle P^{\prime}Oz)$ also denotes the angle between the plane $PO_{l}O_{r}$ and the plane $xOz$, the imaging points $P_{l}$ and $P_{r}$ have the same $\theta$ value in the spherical coordinate. Thus, $P_{l}$ and $P_{r}$ have the same vertical coordinate $u$ on GEER projection images. In other words, the epipolar lines are projected to horizontal lines in GEER domain. As shown in Figure 2(b), although the image structures of projection maps are different for different camera rigs, the matching points in stereo images lie on the same horizontal lines. Therefore, with GEER projection we can transform the two panoramas at arbitrary relative position into left-right stereo pairs that follows the epipolar constraint. Since the matching points have the same $\theta$, the angular disparity $d$ is defined as the difference of $\phi$ : $d=\phi_{l}-\phi_{r}$. The depth of $P$ to the left camera is computed as:
@@ III-B Spherical Sweeping
Inspired by the MVSNet[10] and OmniMVS[7], we utilize the spherical sweeping method to build the unified cost volume with multi-view panoramas. As illustrated in Figure 3, we construct a series of hypothetical spheres at different depths. According to Equation(2), each pixel in the target depth map can be envisioned as representing a ray of light in space ($ray(\theta,\phi)$), and associating it with different depths corresponds to different potential object points along that ray. For each point $P_{i}(\rho_{i},\theta,\phi)$, we can find the corresponding image coordinates of each camera:
@@
In this paper, we introduce two omnidirectional depth estimation mthods that establish geometry constraint based on GEER and Spherical Sweeping method, respectively. We introduce the two algorithms separately in Section IV. Subsequently, we conduct comprehensive experiments to validate and compare the performance of these two methods.
IV Method
@@
IV-A Spherical Feature Extraction Module
Extracting context features from distorted spherical images is challenging for regular CNN modules. In this paper, we implement a Spherical Feature Extraction Module based on spherical convolutions to mitigate the distortion of panoramas. As shown in Figure 4, we implement the sphere convolution based on[61] and accelerate it with CUDA. The sphere convolution changes the sampling pattern to convolve through the neighborhood pixels on the sphere instead of the panorama.
@@ IV-A Spherical Feature Extraction Module
IV-B Pairwise Stereo Matching and Depth Fusion (PSMODE)
@@
IV-B1 Pairwise Stereo Matching
Figure 5 illustrates the process involved in PSMODE. Initially, the multi-view panoramas are organized into multiple stereo image pairs, which are subsequently transformed into GEER projections for pairwise stereo matching. To address distortions, the left and right images are passed through the Spherical Feature Extraction Module to generate feature maps. These feature maps are shifted and concatenated to construct the cost volume. A 3D stacked hourglass network is employed to aggregate the cost volume and estimate the disparity map. The network is optimized with the smoothL1 loss function during training.

### v2

# Robust and Flexible Omnidirectional Depth Estimation with Multiple 360-degree Cameras

@@ Abstract
Omnidirectional depth estimation has received much attention from researchers in 3D perception and measurement in recent years. However, challenges arise due to camera soiling and variations in camera layouts, affecting the robustness and flexibility of the algorithm. In this paper, we use the geometric constraints and redundant information of multiple 360 cameras to achieve robust and flexible multi-view omnidirectional depth estimation. We implement two algorithms, in which the two-stage algorithm obtains initial depth maps by pairwise stereo matching of multiple cameras and fuses the multiple depth maps to achieve the final depth estimation; the one-stage algorithm adopts spherical sweeping based on hypothetical depths to construct a uniform spherical matching cost of the multi-camera images and obtain the depth. Additionally, a generalized epipolar equirectangular projection is introduced to simplify the spherical epipolar constraints. To overcome panorama distortion, a spherical feature extractor is implemented. Furthermore, a synthetic 360 dataset on outdoor road scenes is presented to train and evaluate 360 depth estimation algorithms. Our dataset takes soiled camera lenses and glare into consideration, which is more consistent with the real-world environment. Experiments show that our two algorithms achieve state-of-the-art performance, accurately predicting depth maps even when provided with soiled panorama inputs. The flexibility of the algorithms is experimentally validated in terms of camera layouts and numbers.
keywords:
Omnidirectional Depth Estimation, Omnidirectional 3D Measurement, Spherical Feature Learning, 360Cameras, Autonomous Driving
1 Introduction
Vision-based depth estimation is an essential method for 3D environmental perception. Recently, omnidirectional depth estimation has attracted attention in numerous applications including autonomous driving and robot navigation, owing to its efficiency of the 360 environment. Various algorithms have been proposed to estimate omnidirectional depth maps, including monocular [1, 2, 3], binocular [4, 5] and multi-view approaches [6, 7, 8, 9].
The complex geometric constraints and image distortions of spherical images pose challenges for omnidirectional depth estimation. In addition, the camera may be soiled resulting in image degradation in practical applications. As shown in Fig. 1(e), the images can be soiled by mud spots, water drops or dazzled by intense light. Besides, the camera layouts may vary to accommodate different types of vehicles or robots in real-world tasks. Consequently, the development of an omnidirectional depth estimation algorithm that exhibits robustness against camera soiling and flexibility in adapting to diverse camera configurations becomes imperative and indispensable.
@@ 1 Introduction
Compared to our conference version[11], this extended work encompasses following advancements. Firstly, we expand the applicability of the Cassini projection, to the Generalized Epipolar Equirectangular projection, which accommodates camera pairs at arbitrary relative positions. We provide a thorough analysis and comparison of the spherical geometry constraint models. We introduce the one-stage Spherical-Sweeping MODE and extensively compare its performance with the two-stage Pairwise Stereo matching methods through a wealth of experiments. Furthermore, we demonstrate the flexibility of the proposed methods with varying layouts and numbers of input cameras. Lastly, we present a comprehensive comparative analysis, encompassing the latest state-of-the-art methods, and provide insights for the future advancement of the field.
2 Related Work
2.1 Stereo Matching and Multi-view Stereo Methods
@@
Conventional stereo matching methods estimate disparity map based on the stereo epipolar constraint and image features matching. Some methods aggregate global features to achieve high accuracy, such as SGM[12] and its variants[13, 14, 15], and graph-cut based methods[16, 17]. Deep learning methods report much improved performance in stereo matching. Zbontar and Lecun propose MCCNN[18] that implements the feature extraction with CNNs and computes disparity via conventional cost aggregation. Many methods[19, 20, 21, 22, 23, 24] construct 3D cost volume with image features and optimize the 3D-CNN based cost aggeration modules to estimate disparity maps. Some approaches[25, 26, 27] compute the 2D left-right feature correlation volume. AANet[28] adopts an adaptive aggregation algorithm and replaces the costly 3D-CNNs for an efficient architecture. DMCA-Net[29] utilizes differentiable Markov Random Field for cost aggeration to guide stereo matching. RAFT-Stereo[30] adopts multi-level Gated Recurrent Unit (GRU) to estimate disparity maps recurrently. CREStereo[31] designs a hierarchical network to update disparities iteratively and proposes an adaptive group correlation layer to match points via the local feature.
@@
Multi-view Stereo (MVS) has important applications in 3D reconstruction and has developed rapidly in recent years. Yao et al. [10] proposed the end-to-end MVSNet that builds cost volume by warping feature maps of different views into front-parallel planes of the reference camera to obtain depth maps. P-MVSNet[32] proposes a patch-wise aggregation to build confidence volume and a hybrid network of isotropic and anisotropic 3D-CNNs to exploit context information. Point-MVSNet[33] adopts the feature augmented point cloud to refine the depth map iteratively. Cascade-MVS[34] and CVP-MVS[35] improve the performance with multi-scale coarse-to-fine architectures. UGNet[36] also adopts a coarse-to-fine architecture and leverages uncertainty to improve the depth accuracy. DS-Depth[37] builds the fusion cost volume from multi-frame images to estimate accurate depth maps. PVA-MVSNet[38] proposes self-adaptive view aggregation to generate cost volume instead of the widely-used mean square variance. PVSNet[39] and Vis-MVSNet[40] take the visibility of each view into consideration to supress the mis-matching. Many approaches use the iterative optimization modules to replace the 3DCNNs. R-MVSNet[41] and CER-MVS[42] adopt the GRU module and D2HC-RMVSNet[43] leverages the LSTM module for the cost aggregation. Chen et al.[44] propose a spatial-temporal transformer and leverage self-supervised scheme for multi-view multi-frame depth estimation.
@@
2.2 Omnidirectional Depth Estimation
@@
There are also some methods for obtaining omnidirectional depth maps based on multi-view fisheye cameras. Won et al. propose SweepNet[6] which builds cost volume via spherical sweeping and estimates spherical depth by cost aggregation. They further improve the algorithm and propose the end-to-end OmniMVS[7, 8] architecture to achieve better performance. Meuleman et al. [53] propose an adaptive spherical matching method and an efficient cost aggregation method to achieve real-time omnidirectional MVS. Yang et al. [54] introduce a translation scaling scheme to extend the spherical camera model to multiview for dense 360 depth. OmniVidar[55] adopts the triple sphere camera model and rectifies the multiple fisheye images into stereo pairs of four directions to obtain depth maps. Su et al.[9] leverage a cascade architecture for cost regularization to achieve high accuracy for omnidirectional detph extimation from four fisheye cameras. However, these methods also use planar convolution to extract spherical features and the blind areas of fisheye cameras introduce discontinuity in the spherical cost volume.
2.3 Omnidirectional Depth Datasets
Large-scale datasets with high variety are essential for training and evaluating learning-based algorithms. Recently released omnidirectional depth datasets can be divided into two categories according to the input images, one with the panoramas, and the other with the fisheye images. These datasets are mainly collected from publicly available real-world and synthetic 3D datasets by repurposing them to omnidirectional by rendering. For datasets with panoramas, Wang et al.[49] collect an indoor monocular 360 video dataset named PanoSUNCG from[56]. De La Garanderie et al.[57] provide an outdoor monocular 360 benchmark with 200 images generated from the CARLA autonomous driving simulator[58]. MP3D and SF3D[5] are indoor binocular 360 datasets collected from[59, 60]. 3D60 by Zioulis et al.[48] is an indoor trinocular (central, right, up) 360 dataset collected from[59, 60, 56, 61]. For datasets with fisheye images, Won et al.[6, 7, 8] present three datasets: Urban, OmniHouse and OmniThings. All three datasets are virtually collected in Blender with four fisheye cameras. The fisheye images need complementary information to estimate an omnidirectional depth map, which means discontinuity and requirements for camera directions. In contrast, the panoramas record all 360 information continuously without blind areas. However, as summarized above, the datasets with stereo panoramas consist of indoor scenes only. A detailed summary of multi-view omnidirectional depth datasets can be found in Table 1.
3 Spherical Geometry Constraint Model
@@
To achieve the robust and accurate depth estimation, we establish the geometry constrint of multiple 360 cameras. In this paper, we introduce two spherical geometry constraint models to leverage the multi-view information. In Section 3.1, we introduce the generalized epipolar equirectangular projection, which simplifies the epipolar constraint for binocular panoramas and enables the stereo matching methods on spherical images. In Section 3.2, we present the pipeline of spherical sweeping that builds the cost volume of multi-view panoramas based on the hypothetical sphericals.
3.1 Generalized Epipolar Equirectangular Projection
@@
In this paper, we propose Generalized epipolar equirectangular (GEER) projection to achieve the epipolar constraint for binocular panoramic cameras at arbitrary relative positions in space. As shown in Fig. 2(a), $O_{l}$ and $O_{r}$ are the optic centers of two omnidirectional cameras. We establish a 3D Cartesian coordinate system, where the direction of the x-axis is $O_{r}O_{l}$. $P$ is an object point in 3D space and imaged at points $P_{l}$ and $P_{r}$ on the left and right imaging spheres respectively. $P^{\prime}$ is the projection of $P$ on the plane $yOz$. We define the spherical coordinate system $(\rho,\phi,\theta)$ as follows: $\rho$ is the distance between the object point $P$ and the optic center $O$, $\phi$ is the angle between $PO$ and $x$ axis ($\angle POx$) and denotes the elevation angle, $\theta$ is the angle between $P^{\prime}O$ and $z$ axis on the plane $yOz$ ($\angle P^{\prime}Oz$) and denotes the azimuth angle. Thus, the transformation between Cartesian coordinates and the spherical coordinates is:
@@
where $(u,v)$ denotes the image pixel coordinates in GEER projection and $H,W$ denote the height and width of the image. Because $\theta(\angle P^{\prime}Oz)$ also denotes the angle between the plane $PO_{l}O_{r}$ and the plane $xOz$, the imaging points $P_{l}$ and $P_{r}$ have the same $\theta$ value in the spherical coordinate. Thus, $P_{l}$ and $P_{r}$ have the same vertical coordinate $u$ on GEER projection images. In other words, the epipolar lines are projected to horizontal lines in GEER domain. As shown in Fig. 2(b), although the image structures of projection maps are different for different camera rigs, the matching points in stereo images lie on the same horizontal lines. Therefore, with GEER projection we can transform the two panoramas at arbitrary relative position into left-right stereo pairs that follows the epipolar constraint. Since the matching points have the same $\theta$, the angular disparity $d$ is defined as the difference of $\phi$ : $d=\phi_{l}-\phi_{r}$. The depth of $P$ to the left camera is computed as:
@@ 3.2 Spherical Sweeping
Inspired by the MVSNet[10] and OmniMVS[7], we utilize the spherical sweeping method to build the unified cost volume with multi-view panoramas. As illustrated in Fig. 3, we construct a series of hypothetical spheres at different depths. According to Equation(2), each pixel in the target depth map can be envisioned as representing a ray of light in space ($ray(\theta,\phi)$), and associating it with different depths corresponds to different potential object points along that ray. For each point $P_{i}(\rho_{i},\theta,\phi)$, we can find the corresponding image coordinates of each camera:
@@
In this paper, we introduce two omnidirectional depth estimation mthods that establish geometry constraint based on GEER and Spherical Sweeping method, respectively. We introduce the two algorithms separately in Section 4. Subsequently, we conduct comprehensive experiments to validate and compare the performance of these two methods.
4 Method
@@
4.1 Spherical Feature Extraction Module
Extracting context features from distorted spherical images is challenging for regular CNN modules. In this paper, we implement a Spherical Feature Extraction Module based on spherical convolutions to mitigate the distortion of panoramas. As shown in Fig. 4, we implement the sphere convolution based on[63] and accelerate it with CUDA. The sphere convolution changes the sampling pattern to convolve through the neighborhood pixels on the sphere instead of the panorama.
@@ 4.1 Spherical Feature Extraction Module
4.2 Pairwise Stereo Matching and Depth Fusion (PSMODE)
@@
4.2.1 Pairwise Stereo Matching
Fig. 5 illustrates the process involved in PSMODE. Initially, the multi-view panoramas are organized into multiple stereo image pairs, which are subsequently transformed into GEER projections for pairwise stereo matching. To address distortions, the left and right images are passed through the Spherical Feature Extraction Module to generate feature maps. These feature maps are shifted and concatenated to construct the cost volume. A 3D stacked hourglass network is employed to aggregate the cost volume and estimate the disparity map. The network is optimized with the smoothL1 loss function during training.